f(x) = x^4 + 2x^3 - 252x^2 - 54
Find the interval(s) where f0(x) is decreasing.

Answer by ewatrrr(24785)


Hi
f(x) = x^4 + 2x^3 - 252x^2 - 54
f' = x(4x^2 + 6x - 504) = x(4 - 42)(x + 12)
x = -12, 0 , 10.5
f(x) is decreasing when x < -12 and when 0 < x < 10.5